OpenStudy (anonymous):

How do you write 1-2i in polar form?

OpenStudy (zehanz):

Polar form: a+bi = r(cos(θ)+isinθ). You need r and θ for the polar form. r=√(a²+b²) θ = arctan(b/a) So if you use a=1 and b=-2 in these formulas, you can put them in the polar form r(cos(θ)+isinθ).

OpenStudy (anonymous):

the arctan produces an angle of -63.4. However, when I use the formula x=rcos(theta), I get an angle of +63.4 and 296.6. I'm not sure which angle to use.

OpenStudy (zehanz):

If you put 1-2i in the complex plane, you'll see that is is in the fourth quadrant, so the angle is -63.4.